a paint roller in the shape of a cylinder with a radius of 4cm and a width of 21cm is rolled vertically on a wall.
a) What is the wet path after ten compete rolls
b) what area does the paint cover

Answer by mananth(16946)

a paint roller in the shape of a cylinder with a radius of 4cm and a width of 21cm is rolled vertically on a wall.
a) What is the wet path after ten compete rolls
b) what area does the paint cover
If it rolls once the length it rolls is one circumference=2pir
=2*pi*4
=8pi
after 10 rolls it will roll a length of 80 pi=251.2 cm
The wet path is 21 cm by 251.2 cm
area = 21*251.2=5275.2 cm^2